# Magic-Boost: Boost 3D Generation with Mutli-View Conditioned Diffusion

### v1

@@ Abstract
Benefiting from the rapid development of 2D diffusion models, 3D content creation has made significant progress recently. One promising solution involves the fine-tuning of pre-trained 2D diffusion models to harness their capacity for producing multi-view images, which are then lifted into accurate 3D models via methods like fast-NeRFs or large reconstruction models. However, as inconsistency still exists and limited generated resolution, the generation results of such methods still lack intricate textures and complex geometries. To solve this problem, we propose Magic-Boost, a multi-view conditioned diffusion model that significantly refines coarse generative results through a brief period of SDS optimization ($\sim 15$min). Compared to the previous text or single image based diffusion models, Magic-Boost exhibits a robust capability to generate images with high consistency from pseudo synthesized multi-view images. It provides precise SDS guidance that well aligns with the identity of the input images, enriching the local detail in both geometry and texture of the initial generative results. Extensive experiments show Magic-Boost greatly enhances the coarse inputs and generates high-quality 3D assets with rich geometric and textural details project page.
@@
Figure 1: Provided with an input image and its coarse 3D generation, MagicBoost effectively boosts it to a high-quality 3D asset within 15 minutes. From left to right, we show the input image, pesudo multi-view images and coarse 3D results from Instant3D, together with the significantly improved results produced by our method.
1 Introduction
The recent surge in the development of 2D diffusion models has unlocked new prospects for 3D content generation. A particularly viable strategy involves fine-tuning pre-trained 2D diffusion models to catalyze their capacity for generating images with multi-view consistency, which are then lifted to accurate 3D models through fast-NeRFs or large reconstruction models [17, 19, 13, 20]. For example, Instant3D [13] firstly finetune the pre-trained 2D diffusion models to unlock the ability of multi-view image generation, and then utilize a robust reconstruction model to derive 3D representations; Wonder3D [19] finetunes the 2D diffusion model with cross-domain attention layers to enhance the 3D consistency of generative outputs. Despite the efficiency of these methods in producing 3D assets from textual prompts or images, the generated results are still coarse, characterized by a lack of fine textures and complex geometries. This is primarily caused by local inconsistencies and the limited resolution of the generative process.
@@ 1 Introduction
A potential solution is to enhance the generated results of such methods with Score Distillation Sampling (SDS) optimization [18, 5]. Commencing with a coarse 3D model, efforts have been made to refine it through SDS optimization with small noise levels, utilizing text or single-view conditioned diffusion models [16, 26]. However, we argue that both text and single-view image conditions are inadequate in providing explicit control and precise guidance. This inadequacy often results in identity shifts, blurred textures, and geometric inaccuracies. Specifically, the inherent ambiguity of textual descriptions poses a challenge for text-conditioned diffusion models, such as StableDiffusion and DeepFoldy-IF, to maintain a consistent identity throughout the optimization process. Consequently, the optimized results may diverge significantly from the initial models, potentially defying users’ expectations. Single-view image-guided diffusion models like Zero-1-to-3 [16] endow 2D diffusion models with the capacity for viewpoint conditioning. Nevertheless, constrained by the limited information provided bt single-view inputs, Zero-1-to-3 and subsequent methods [16, 30, 17, 19, 3, 41] are prone to generating unpredictable and implausible shapes and textures during novel view prediction or SDS optimization, leading to flattened geometries and blurry textures.
To tackle the challenges in refining coarse 3D generations, we present Magic-Boost, a multi-view conditioned diffusion model that ingests multi-view images as inputs, implicitly encoding 3D information across different views. Magic-Boost exhibits a robust capability to generate images with high 3D consistency as well as provide precise SDS guidance that aligns well with the identity of the input images, enriching the local detail in both geometry and texture of the initial generated results. Built upon the Stable Diffusion architecture, we introduce an advanced method to efficiently distill dense local features from multi-view inputs with a denoising U-Net operating at a fixed time step. In line with prior arts [31, 19, 34], our model employs a self-attention mechanism to enable interactions and information sharing across different views, thus implicitly encoding the multi-view correlations. However, relying exclusively on ground-truth multi-view images for training produces unsatisfying results. We meticulously develop a series of data augmentation strategies, including random drop, random scale, and noise disturbance, to facilitate the training process, leading to more robust performance. We also introduce a condition label that allows users to manually adjust the influence of different input views, strength the influence of high-fidelity input views while mitigating that of lower-quality ones. To further enhance the optimization process, we present a novel Anchor Iterative Update loss to address the over-saturation problem in SDS, culminating in the generation of high-quality content with realistic textures.
Figure 2: The overall pipeline. We build MagicBoost with Instant3D [13] (a synergy of Multi-view Diffusion (MVDiff) and Large Reconstruction Model (LRM)), which provides pseudo multi-view images together with a coarse 3D output. Conditioned on the generated multi-view inputs, our model provides precise SDS guidance, thereby significantly augmenting the coarse 3D outputs within a brief interval ($\sim 15$min).
As shown in Fig. 2, our testing framework integrates Instant3D [13], a two-stage feed-forward generative approach, which firstly generates four multi-view images with finetuned 2D diffusion models, followed by the utilization of a large reconstruction model for 3D reconstruction. Given the four synthesized multi-view images from Instant3D as inputs, our model is capable of synthesizing highly consistent novel views and provides precise SDS guidance, refining the coarse generated results within a short time period ($\sim 15$min). Comprehensive evaluations demonstrate Magic-Boost substantially enhances the quality of coarse inputs, efficiently yielding 3D assets of better quality with intricate geometries and authentic textures, as shown in Fig. 1.
2 Related Works
@@
3D generation Models. Traditional GAN-based methods explore generation of 3D models with different 3D representations including voxels [10, 32], point clouds [38, 1] and meshes [7, 37]. These methods, predominantly trained on limited-scale 3D datasets, often fall short in generating intricate geometric structures with substantial diversity. The advent of diffusion models in 2D generative tasks has prompted several explorations into their application within 3D domains, training diffusion models directly on 3D datasets. For example, Shap-E [11] trains an conditional diffusion models by encoding each 3D shape into a set of parameters of an implicit function. Nonetheless, the generative quality of such methods remains constrained by the scarce availability and variety of 3D datasets. Even though the largest 3D dataset [4] in recent years are still much smaller than the datasets used for 2D image generation training [29]. Recently, a novel paradigm has emerged for 3D generation that circumvents the need for large-scale 3D datasets by leveraging pretrained 2D generative models. Leveraging the semantic understanding and high-quality generation capabilities of pretrained 2D diffusion model, Dreamfusion [25], for the first time, propose to optimize 3D representations directly with 2D diffusion model utilizing the Score distilling Sampling loss. Follwoing works [26, 15, 39, 2] continue to enhance various aspects such as generation fidelity and training stability by proposing different variations of SDS such as VSD [35]. Although being able to generate high-quality 3D contents, these methods suffer from extremely long time for optimization, which greatly hinders its application in the real-world scenario. Another line of works try to finetune the pretrained 2D diffusion network to unlock its ability of generating multi-view images simultaneously, which are subsequently lifted to 3D models with fast-NeRF [24] or large reconstruction models [17, 19, 13, 20]. Although these methods generates reasonable results, they are still limited by the local inconsistency and limited generation resolution, producing coarse results without detailed textures and complicated geometries.
@@
Novel View Synthesis. The success of diffusion model on the 2D task has opened a new door for the task of zero-shot novel view synthesis. Finetuned on Stable Diffusion, Zero-1-to-3 [16] achieves viewpoint-conditioned image synthesis of an object from a single in-the-wild. Subsequent [41, 30, 17] works further improves the generative quality by refining network architectures and optimizing training methodologies. Yet, constrained by the limited scope of single-view input, these models encounter challenges in producing accurate novel views with robust 3D consistency. In contrast, EscherNet [12] propose to learn implicit and generative 3D representations coupled with specialised camera positional encodings, allowing the network to fuse information from arbitrary number of reference and target views, which greatly improves the 3D understanding of the network, leading to better generation results. In parallel with EscherNet, our model also employs multi-view inputs to enhance the accuracy of the 3D generation. However, we operate within a more challenging task—enhancing the local details of the coarse 3D model with generated pseudo multi-view images as inputs.
3 Methods
@@
Given a coarse 3D asset and its corresponding multi-view generated pseudo images, we aim to improve the asset’s quality, enriching both geometry and texture through SDS optimization within a short time period. To achieve this, we propose Magic-Boost, a diffusion model conditioned on multiple views that adeptly encodes 3D information and ensures multi-view consistency. This model is designed to provide precise SDS guidance, thereby refining the local details of the initial coarse results. Magic-Boost is built upon the Stable Diffusion architecture, where we adeptly extract dense local features via a denoising U-Net operating at a fixed timestep and adopt the self-attention mechanism to facilitate interactions and information exchange across differernt views, as detailed in Sec. 3.1. To enhance the training process, we introduce several data augmentation strategies, as elaborated in Sec. 3.2. Our testing framework is built on Instant3D. In the refinement phase, we introduce an Anchor Iterative Update loss to alleviate the over-saturation problem of SDS, leading to high-quality generation results with detailed geometry and realistic textures, as presented in Sec. 3.3.
@@
Figure 3: Architecture of our multi-view conditioned diffusion model. At the core of our model lies the extraction of dense local features facilitated by a denoising U-Net operating at a fixed timestep. Concurrently, we harness a frozen CLIP ViT encoder to distill high-level signals. The original 2D self-attention layer is extended into 3D by concatenating keys and values across various views. To further control the condition strength of different views, we involve a condition label which allows users to manually control the condition strength of different input views independently.
3.1 Multi-view Conditioned Diffusion
@@
Formulation. Given $n$ views $v_{0},v_{1},...,v_{n}\in\mathbb{R}^{H\times W\times 3}$, each capturing an object from a distinct perspective with relative angles $\Delta\gamma_{0}=0,\Delta\gamma_{1},...,\Delta\gamma_{n}$ of an object as input, the objective of our multi-view conditioned diffusion model is to synthesize a novel view $x$ at a relative angle $\Delta\gamma_{x}$. To this end, we compute the camera rotation $R_{i}\in\mathbb{R}^{3\times 3}$ and translation $T_{i}\in\mathbb{R}^{3}$ corresponding to each relative angle $\Delta\gamma_{i},i\in(0,1,...,n,x)$ and subsequently train the model $\mathcal{M}$ to generate the novel view $x$. The relative camera pose is denoted as $c_{i}=(R_{i},T_{i}),i\in(0,1,...,n,x)$, and the formulation of our model is as follows:
$x_{c_{x}}=\mathcal{M}(v_{0},...,v_{n},c_{0},...,c_{n},c_{x})$ (1)
@@
In our experiments, we follow the setting of Instant3D and employing a four-view arrangement, where the four condition views are orthogonal to each other with relative angles set at $\Delta\gamma_{0}=0,\Delta\gamma_{1}=90,\Delta\gamma_{2}=180$, and $\Delta\gamma_{3}=270$.
@@
Local Feature Extractor. Multi-view images provide dense local details, pivotal for novel view synthesis. However, accurately and efficiently encoding local dense features from multi-view inputs is non-trivial. Zero-1-to-3 [16] attempts to encode dense local signals by appending the reference image to the input of the denoising U-Net within Stable Diffusion. This approach, however, enforces an misaligned pixel-wise spatial correspondence between the input and target images. Zero123++ [30] extracts local feature by processing an additional reference image through the denoising U-Net model. It synchronizes the Gaussian noise level of the input images with that of the denoising input, enabling the U-Net to focus on pertinent features at varying noise levels. While this method is adept at extracting dense local features, it leads to much more computational costs, by extracting different feature maps at each denoising step. To address this issue, we propose a novel technique that extracts dense local features using the denoising U-Net at a fixed timestep. Specifically, we employ the denoising U-Net to extract low-level features from clean images without introducing any noise, and we consistently set the timestep to zero. This approach not only captures features with dense local details but also significantly accelerates the generation process by extracting local features just once throughout the entire optimization or diffusion procedure.
@@
Multi-view Conditioned Generation. The overall architecture of our model is shown in Fig. 3. Our multi-view conditioned diffusion model is built on Stable Diffusion backbone. We incorporate global features via the cross-attention module, following methods such as IP-Adapter [40] and ImageDream [34]. The denoising U-Net, operating at a fixed timestep, is harnessed to distill dense local features from multi-view inputs. These features are subsequently integrated into the self-attention module to encode 3D correspondence. Similar with prior arts [31, 19], we extend the original 2D self-attention layer of Stable Diffusion into 3D by concatenating the keys and values of different views within the self-attention layers, which facilitates the interactions across different input views and implicitly encodes multi-view correspondence. Additionally, camera poses are injected into the denoisinge U-Net for both the conditional multi-view inputs and the target view. Specifically, a two-layer MLP is employed to encode camera poses into one-dimensional embeddings, which are then added to time embeddings as residuals.
3.2 Data Augmentation
@@
We train the multi-view conditioned diffusion model with ground-truth multi-view images rendered from Objaverse [4]. However, directly training with these ground-truth images can lead to suboptimal results during inference, as the domain discrepancy between the ground-truth multi-view images and the synthesized ones used during testing engenders artifacts and inconsistent generative results. Moreover, when trained with multi-view images, the model may develop a bias towards the nearest anchor view, neglecting information from other views, which can result in inaccurate outputs that fail to maintain 3D consistency
To solve these problems, we propose several data augmentations strategies to enhance the training process and ensuring robust performance:
  • Noise Disturb and Random Scale: To emulate the blurry textures and local inconsistencies present in synthesized multi-view images, we introduce noise disturb augmentation, which perturbs the conditional multi-view images with Gaussian noise sampled at different timesteps. This is coupled with random scale augmentation, employing a downsample-and-upsample approach to generate blurry training inputs, thereby enhancing the model’s capability to handle such imperfections.
@@
  • Random Drop: To circumvent the model’s tendency to rely excessively on the nearest conditional view, we implement random drop augmentation. This technique randomly omits conditional views during training, enforcing the network to synthesize more consistent results by integrating information from all available views.
@@
Similar as LGM [33], we also incorporate grid distortion to disrupt the 3D coherence of ground-truth images and camera jitter augmentation to vary the conditional camera poses of multi-view inputs.
@@
To further control the influence of different views, we introduce a condition label enabling manual adjustment of the input views’ conditioning strength. This label, indicative of the conditional intensity, is processed by a two-layer MLP into a one-dimensional vector, which is then combined with the time embedding in the local feature extractor to guide the generation process. For the target view, the condition label is consistently zero. This condition label is concurrently trained with our data augmentation strategy. When employing more augmentation scales, such as noise perturbation with larger timesteps, we assign a higher value to the condition label to refuce the condition weights, and conversely for smaller scales.
3.3 Refinement with SDS Optimization
Figure 4: Illustration of the anchor iterative update loss.

### v3

# Magic-Boost: Boost 3D Generation with Multi-View Conditioned Diffusion

@@ Abstract
Benefiting from the rapid development of 2D diffusion models, 3D content generation has witnessed significant progress. One promising solution is to finetune the pre-trained 2D diffusion models to produce multi-view images and then reconstruct them into 3D assets via feed-forward sparse-view reconstruction models. However, limited by the 3D inconsistency in the generated multi-view images and the low reconstruction resolution of the feed-forward reconstruction models, the generated 3d assets are still limited to incorrect geometries and blurry textures. To address this problem, we present a multi-view based refine method, named Magic-Boost, to further refine the generation results. In detail, we first propose a novel multi-view conditioned diffusion model which extracts 3d prior from the synthesized multi-view images to synthesize high-fidelity novel view images and then introduce a novel iterative-update strategy to adopt it to provide precise guidance to refine the coarse generated results through a fast optimization process. Conditioned on the strong 3d priors extracted from the synthesized multi-view images, Magic-Boost is capable of providing precise optimization guidance that well aligns with the coarse generated 3D assets, enriching the local detail in both geometry and texture within a short time ($\sim 15$min). Extensive experiments show Magic-Boost greatly enhances the coarse generated inputs, generates high-quality 3D assets with rich geometric and textural details.
Index Terms:
3D Generation, Multi-view Diffusion, Image to 3D Generation
@@
Figure 1: Provided with an input image and its coarse 3D generation, MagicBoost effectively boosts it to a high-quality 3D asset within 15 minutes. From left to right, we show the input image, pesudo multi-view images and coarse 3D results from Instant3D [1], together with the significantly improved results produced by our method.
I Introduction
The recent surge in the development of 2D diffusion models has opened a new door for 3D content generation. One of the promising methods is to first synthesize multi-view consistent images with fine-tuned 2D diffusion model and then reconstruct them into 3D assets through fast-NeRFs [2, 3, 4] or large reconstruction models [1, 5]. Despite the efficiency of these methods, the generated results still suffer from artifacts like coarse textures or incorrect geometries, which is primarily caused by the local inconsistencies existing in the synthesized multi-view images and limited reconstruction resolution from the feed-forward reconstruction models.
@@ I Introduction
Recently, efforts [6, 7, 8] have been made to adopt SDS (Score Distillation Sampling [9]) to further refine the coarse generated results through an optimization process. However, SDS optimization process is hard to control and leads to instable refinement results with problems like identity shifts, blurred textures, and geometric collapse [9, 10, 6]. We argue that the main reason for this instability is the lack of 3D understanding and explicit identity control ability of 2D diffusion models on the optimization process. Specifically, the inherent ambiguity of textual descriptions poses a challenge for text-conditioned diffusion models, such as StableDiffusion [11] and DeepFoldy-IF [12], to maintain a consistent identity throughout the optimization process. As a result, the refined results may diverge significantly from the initial coarse 3D assets, which potentially defies users’ expectations. On the other hand, single-view image conditioned diffusion models like Zero-1-to-3 [13] empower 2D diffusion models with the view-condition capacity. However, the information provided by single-view images is very limited, leading to implausible shapes and blurry textures of the refined results.
To address this problem, we introduce a multi-view based refine method, named Magic-Boost, to further refine the generation results. The main motivation of us is that the synthesized multi-view images, despite not strictly 3D consistent, are still capable of providing strong 3D priors which align well with the target generation results and could be extracted by the multi-view conditioned diffusion model to guide the refinement process. Inspired by this, we first introduce a novel multi-view conditioned diffusion model that adopts synthesized pseudo multi-view images as inputs, implicitly distill 3D information across different views to synthesize high-fidelity novel view images and provide precise SDS (Score Distillation Sampling [9]) guidance to refine the local details of the coarse generated 3D assets. Specifically, we build the model upon the Stable Diffusion architecture [11], further equip it with a novel time-fixed local feature extractor to efficiently capture the low-level local details and enable the interactions across different views with cross-frame 3D attention. During training stage, we meticulously develop a series of data augmentation strategies to imitate the 3D inconsistency to empower the model to extract strong 3D priors from inconsistent multi-view inputs at test time. During the optimization process, we present a novel Anchor Iterative Update loss to address the over-saturation problem in the Score Distillation Sampling process [9], culminating in the generation of high-quality content with realistic textures.
Figure 2 depicts the overall pipeline of the proposed Magic-Boost refinement method. Magic-Boost could be seen as a plug-in module which could be pluged into any 3D generation methods capable of providing pseudo multi-view priors, such as Instant3D [1], InstantMesh [14], LGM [15] and etc. As shown in Figure 1, benefit from the strong 3d priors provided by the pesudo multi-view images, Magic-boost provides precise SDS guidance, significantly enhancing the coarse 3D outputs within a brief interval ($\sim 15$min). Comprehensive evaluations demonstrate Magic-Boost substantially enhances the quality of coarse inputs, efficiently yielding 3D assets of better quality with intricate geometries and authentic textures.
Figure 2: The overall pipeline. The proposed Magic-Boost could be a plug-in module plugged into any 3D generation methods capable of providing pseudo multi-view priors, such as Instant3D [1], InstantMesh [14], LGM [15] and etc. Benefit from the strong 3d priors provided by the pesudo multi-view images, Magic-boost provides precise SDS guidance, significantly enhancing the coarse 3D outputs within a brief interval ($\sim 15$min).
II Related Works
@@
3D generation Models. Traditional GAN-based methods explore generation of 3D models with different 3D representations including voxels [16, 17], point clouds [18, 19, 20, 21] and meshes [22, 23, 24, 25]. Recently, the advance of diffusion models in 2D generative tasks [11, 12, 26] has prompted explorations of their application into 3D domains. Efforts [27] have been made to directly train diffusion models on 3D datasets. However, even though the largest 3D dataset [28] in recent years are still much smaller than the datasets used for 2D image generation training [29]. Recently, a novel paradigm has emerged for 3D generation that circumvents the need for large-scale 3D datasets by leveraging pretrained 2D generative models. Leveraging the semantic understanding and high-quality generation capabilities of pretrained 2D diffusion model, Dreamfusion [9], for the first time, propose to optimize 3D representations directly with 2D diffusion model utilizing the Score distilling Sampling loss. Follwoing works [8, 30, 31, 32] continue to enhance various aspects such as generation fidelity and training stability by proposing different variations of SDS such as VSD [10]. Although being able to generate high-quality 3D contents, these methods suffer from extremely long time for optimization, which greatly hinders its application in the real-world scenario. Another line of works try to finetune the pretrained 2D diffusion network to unlock its ability of generating multi-view images simultaneously, which are subsequently lifted to 3D models with fast-NeRF [33] or large reconstruction models [2, 3, 1, 5]. Although these methods generates reasonable results, they are still limited by the local inconsistency and limited generation resolution, producing coarse results without detailed textures and complicated geometries.
@@
Novel View Synthesis. The success of diffusion model on the 2D task has opened a new door for the task of zero-shot novel view synthesis. Finetuned on Stable Diffusion [11], Zero-1-to-3 [13] achieves viewpoint-conditioned image synthesis ability on different inputs. Subsequent [34, 4, 2] works further improve the generative quality of Zero-1-to-3 by refining the network architecture and improving the training methodologies. Yet, constrained by the limited information provided by single-view input, these models encounter challenges in producing accurate novel views with robust 3D consistency. In contrast, EscherNet [35] propose to improve the 3D understanding of the network by fusing information from arbitrary number of reference and target views, leading to better view synthesis and reconstruction results. Compared to EscherNet, our model focus on a more challenging generation task: employing pseudo multi-view images as 3d priors to enhance the accuracy of the 3D generation.
III Methods
@@
We propose Magic-Boost, a multi-view conditioned diffusion model that explores the strong 3d prior provided by the synthesized multi-view images to refine the coarse generated results. The proposed model is built upon the Stable Diffusion architecture, where we extract dense local features via a denoising U-Net operating at a fixed timestep and adopt the self-attention mechanism to achieve information exchange across different views, as detailed in Sec. III-A. To empower the model to extract strong 3D priors from inconsistent multi- view inputs at test time, we introduce several data augmentation strategies, as elaborated in Sec. III-B. In the refinement phase, we introduce an Anchor Iterative Update loss to alleviate the over-saturation problem of SDS [9], as presented in Sec. III-C.
@@
Figure 3: Architecture of our multi-view conditioned diffusion model. At the core of our model lies the extraction of dense local features facilitated by a denoising U-Net operating at a fixed timestep. Concurrently, we harness a frozen CLIP ViT encoder to distill high-level signals. The original 2D self-attention layer is extended into 3D by concatenating keys and values across various views. To further control the condition strength of different views, we involve a control label which allows users to manually control the condition strength of different conditional views.
III-A Multi-view Conditioned Diffusion
@@
Formulation. Given $n$ views $v_{0},v_{1},...,v_{n}\in\mathbb{R}^{H\times W\times 3}$, each capturing an object from a distinct perspective with relative angles $\Delta\gamma_{0}=0,\Delta\gamma_{1},...,\Delta\gamma_{n}$ as input, the objective of our multi-view conditioned diffusion model is to synthesize a novel view $x$ at a relative angle $\Delta\gamma_{x}$. To this end, we compute the camera rotation $R_{i}\in\mathbb{R}^{3\times 3}$ and translation $T_{i}\in\mathbb{R}^{3}$ corresponding to each relative angle $\Delta\gamma_{i},i\in(0,1,...,n,x)$ and subsequently train the model $\mathcal{M}$ to generate the novel view $x$. The relative camera pose is denoted as $c_{i}=(R_{i},T_{i}),i\in(0,1,...,n,x)$, and the formulation of our model is as follows:
$x_{c_{x}}=\mathcal{M}(v_{0},...,v_{n},c_{0},...,c_{n},c_{x})$
@@
In our experiments, we follow the setting of Instant3D [1] and employing a four-view arrangement, where the four condition views are orthogonal to each other with relative angles set at $\Delta\gamma_{0}=0,\Delta\gamma_{1}=90,\Delta\gamma_{2}=180$, and $\Delta\gamma_{3}=270$.
@@
Local Feature Extractor. Multi-view images provide dense local details, pivotal for novel view synthesis. However, accurately and efficiently encoding local dense features from multi-view inputs is non-trivial. Zero-1-to-3 [13] attempts to encode dense local signals by appending the reference image to the input of the denoising U-Net within Stable Diffusion. This approach, however, enforces an misaligned pixel-wise spatial correspondence between the input and target images. Zero123++ [4] extracts local feature by processing an additional reference image through the denoising U-Net model. It synchronizes the Gaussian noise level of the input images with that of the denoising input, enabling the U-Net to focus on pertinent features at varying noise levels. While this method is adept at extracting dense local features, it leads to much more computational costs, by extracting different feature maps at each denoising step. To address this issue, we propose a novel technique that extracts dense local features using the denoising U-Net at a fixed timestep. Specifically, we employ the denoising U-Net to extract low-level features from clean images without introducing any noise, and we consistently set the timestep to zero. This approach not only captures features with dense local details but also significantly accelerates the generation process by extracting local features just once throughout the entire SDS optimization or diffusion procedure.
@@
Multi-view Conditioned Generation. The overall architecture of our model is shown in Figure 3. Our multi-view conditioned diffusion model is built on Stable Diffusion backbone [11]. We incorporate global features via the cross-attention module, following methods such as IP-Adapter [36] and ImageDream [38]. The denoising U-Net, operating at a fixed timestep, is utilized to distill dense local features from multi-view inputs. These features are subsequently integrated into the self-attention module to encode 3D correspondence. Similar with [40, 3], we extend the original 2D self-attention layer of Stable Diffusion into 3D by concatenating the keys and values of different views within the self-attention layers, which facilitates the interactions across different input views and implicitly encodes multi-view correspondence. Additionally, camera poses are injected into the denoisinge U-Net for both the conditional multi-view inputs and the target view. Specifically, a two-layer MLP is employed to encode camera poses into one-dimensional embeddings, which are then added to time embeddings as residuals.
III-B Data Augmentation
@@
We train the multi-view conditioned diffusion model with ground-truth multi-view images rendered from Objaverse [28] as condition. However, directly training with these ground-truth images can lead to suboptimal results during inference, as the domain discrepancy between the ground-truth multi-view images and the synthesized ones used during testing leads to artifacts and inconsistent generation results. To empower the model to extract strong 3D priors from inconsistent multi-view inputs at test time, we introduce several data augmentation strategies during the training stage:
  • Noise Disturb and Random Scale: To imitate the blurry textures and local inconsistencies present in synthesized multi-view images, we introduce noise disturb augmentation. Specifically, we perturb the clean conditional multi-view images with random Gaussian noises, which is similar as the forward diffuse process. During our experiments, we adopt the DDPM forward strategy and uniformly sample noises under $t=300$ to destroy the conditional mutli-view images, which are then used as conditional inputs to train the multi-view conditioned diffusion model. To further enhance the networks’ robustness towards blurry conditional inputs, we propose a random scale augmentation strategy, which employs a downsample-and-upsample approach to generate blurry training inputs.
@@
  • Random Drop: During training, we observed that the model tends to learn strong bias by relying heavily on the views closest to the target view for generating novel views, while ignoring the influence of other conditional views. This leads to artifacts and instable results as errors may exist in the closest conditional views. To eliminate this tendency, we introduce random drop augmentation, by randomly dropping different conditional views during training, enforcing the network to synthesize more consistent results by integrating information from all available views.
@@
Similar as LGM [15], we also incorporate grid distortion to disrupt the 3D coherence of ground-truth images and camera jitter to vary the conditional camera poses of multi-view inputs.
@@
To simulate the test scenario, we set the first sampled view as the default input image (simulating user inputs at test stage) and only apply data augmentation to the other conditional views. To further control the influence of different views, we introduce a control label enabling manual adjustment of the conditioning strength for different conditional views. This label, indicative of the conditional strength, is processed by a two-layer MLP into a one-dimensional vector, which is then combined with the time embedding in the local feature extractor to guide the generation process. During training stage, When employing larger augmentation scales, such as noise perturbation with larger timesteps, we assign a lower value to the condition label to reduce the control weights, and vice versa.
Figure 4: Illustration of the anchor iterative update loss. In detail, we regard the input pesudo multi-view inputs as our initial anchor datasets and adopt an update strategy by first rendering anchor view image, perturbing the image with random noise and then apply a multi-step denoising process with the proposed multi-view condition diffusion model to refine the anchor images. The refined anchor images are then used to supervise the generation with MSE loss to eliminate the over-stature problem during the SDS optimization process [9].
III-C Refinement with SDS Optimization